# SeagrassFinder: Deep Learning for Eelgrass Detection and Coverage Estimation in the Wild

### v1

@@ Abstract
Seagrass meadows play a crucial role in marine ecosystems, providing important services such as carbon sequestration, water quality improvement, and habitat provision. Monitoring the distribution and abundance of seagrass is essential for environmental impact assessments and conservation efforts. However, the current manual methods of analyzing underwater video transects to assess seagrass coverage are time-consuming and subjective. This work explores the use of deep learning models to automate the process of seagrass detection and coverage estimation from underwater video data. A dataset of over 8,300 annotated underwater images was created, and several deep learning architectures, including ResNet, InceptionNetV3, DenseNet, and Vision Transformer, were evaluated for the task of binary classification of “Eelgrass Present” and “Eelgrass Absent” images. The results demonstrate that deep learning models, particularly the Vision Transformer, can achieve high performance in predicting eelgrass presence, with AUROC scores exceeding 0.95 on the final test dataset. The use of transfer learning and the application of the Deep WaveNet underwater image enhancement model further improved the models’ capabilities. The proposed methodology allows for the efficient processing of large volumes of video data, enabling the acquisition of much more detailed information on seagrass distributions compared to current manual methods. This information is crucial for environmental impact assessments and monitoring programs, as seagrasses are important indicators of coastal ecosystem health. Overall, this project demonstrates the value that deep learning can bring to the field of marine ecology and environmental monitoring.
Index Terms:
Deep Learning, Marine Biology, Marine Ecology, Marine Imaging, Computer Vision, Transfer Learning
I Introduction
Seagrasses are a type of marine flowering plant that are of fundamental importance in the health and integrity of coastal ocean habitats, biodiversity, global climate, human economic stability (Cullen-Unsworth et al., 2014) as well as global food security (Unsworth et al., 2019). Despite ongoing conservation efforts, seagrass meadows are declining globally at an alarming rate due to human activities and coastal development, and exacerbated further by climate change (Waycott et al., 2009). These factors collectively contribute a stark decline water quality and threaten the long-term survival of seagrass habitats. There is a growing interest in offshore infrastructure such as windfarms which offer a potential solution to the energy crisis but also raise concerns about environmental impacts. We need to effectively measure and quantify the impacts of these interventions to the local underwater environments. Environmental Impact Assessments (EIAs) conducted by private companies are designed to oversee offshore building sites and their impacts to ensure environmental compliance with different stakeholders across the marine infrastructure sector. Monitoring is often conducted in form of video transects; a camera is mounted onto a sled and towed behind a vessel to capture underwater video footage. The resulting recorded videos are then meticulously annotated by marine biologists.
One crucial challenge here is data acquisition and repeated monitoring of the condition and distribution of such seagrass meadows in a structured, automated and efficient manner (Unsworth et al., 2019). Some automation approaches like remote sensing have shown success but fail to distinguish between different types of seagrass (Riegl and Purkis, 2005). Recently, there is growing interest in detecting and classifying seagrasses directly from underwater (UW) image or video data taken in the wild to gain more fine-grained knowledge on the health of seagrass meadows. Acquiring this visual data is very costly in terms of human efforts and operational investments as it requires expertise not only in technical equipment but also in marine biology. It is a priority to overcome these costs and challenges by automation, specifically by applying image analysis.
With the advent of deep learning, our capacity to analyze and process data has transformed significantly. Specifically deep neural networks (DNNs) have shown remarkable success in a wide range of computer vision tasks such as image classification, semantic segmentation and object detection. DNNs achieve this by learning information directly from raw pixel input data without the need for manual feature engineering. Especially convolutional neural networks (Krizhevsky et al., 2012) massively improved our ability to automatically extract relevant features and have become the new standard in computer vision tasks. Successfully training DNNs with high accuracy is dependent on the quality and quantity of input data, the training procedure and the type of DNN models used. The advantage here is that this allows for adaptation to a variety of image qualities and conditions (such as blurry, color-distorted or underexposed underwater images) as long as labeled samples are sufficiently present in the training set. However this also marks a challenge since DNNs require large amounts of such samples which in turn requires vast human annotation efforts.
The objective of this work is to address all above-mentioned limiting factors of training DNNs for seagrass detection in the wild by; (1) deploying an efficient image annotation platform to gather large amounts of labeled data (2) implementing an effective training procedure for our underwater video data set, (3) investigating a range of DNN models on their capacity to learn seagrass features via transfer learning. The data used in this project was taken from DHI’s Environmental Impact Assessment (EIA) and baseline reporting of the Lynetteholm project, a new artificial peninsula being created to extend the coastline of northern Copenhagen in Denmark.
@@ I Introduction
Figure 1: The SeagrassFinder Project Pipeline: a vessel tows a sled with a camera along the seabed and records videos of the UW environment. The extracted images are labeled by human annotators and used to train a DNN on the task of detecting eelgrass. The trained DNN can now be used to replace the time-consuming manual annotation process.
@@
We address the challenge of estimating seagrass presence and coverage in the wild from vast amounts of raw, unstructured and unfiltered video data taken in realistic environmental conditions. We achieve this by deploying a custom-made annotation platform for simplified manual labeling, as well as training state-of-the-art deep learning models with a specific focus on classifying whether eelgrass is present or absent in an image. Specifically we contribute the following:
@@
    a strategy for effectively streamlining human labeling efforts of experts in the field as well as non-experts on underwater seagrass images and compiling the annotated data set into an efficient ML data pipeline,
@@
    an analysis of state-of-the-art DNN models and their ability to classify a range of challenging underwater images according to the presence eelgrass using a transfer learning mechanism, as well as an investigation of the effect of UW image enhancement on DNN performance,
@@
    a new experimental approach for further utilizing above-mentioned DNN models for estimating visual eelgrass coverage for an image stream from binary predictions as a post-processing step,
@@
    a pipeline for significantly reducing the time and financial resources required for human annotation tasks and paving the way towards potential automated environmental monitoring of seagrass for example by means of autonomous underwater vehicles.
To the best of our knowledge we are the first to deploy state-of-the art Vision Transformers to detect the presence of eelgrass in images in this line of research.
I-A Seagrass
Seagrasses are a type of submerged aquatic vegetation (SAV) found in shallow waters along every continent in the world except Antarctica (Novak and Short, 2020). Seagrass meadows are monitored worldwide as part of environmental impact assessments (EIAs) and environmental monitoring programs because they provide important ecosystem services, such as carbon sequestration, water quality improvement, providing food and habitat, and acting as a biological indicator of local ecosystems (Novak and Short, 2020). Seagrass meadows play a crucial role as a major carbon sink worldwide and are essential for mitigating the impacts of global climate change. Although seagrasses cover only 2% of the ocean’s seabed, they account for over 15% of the carbon annually sequestered in marine environments (Fourqurean et al., 2012). Besides their contributions to carbon sequestration and exportation, seagrass meadows also act as filters, enhancing water quality and clarity by directly capturing suspended particles and retaining organic matter (Terrados and Duarte, 2000). Seagrass meadows serve as vital hubs of biodiversity, offering nourishment and shelter to a wide range of organisms, including endangered species like dugongs and commercially significant ones such as fish and shrimp, alongside microbes and invertebrates (Fry and Parker, 1979). Furthermore, due to seagrasses significant vulnerability and rapid response to anthropogenic stressors such as eutrophication, sedimentation, oil spills and commercial fishing, seagrasses also act as biological indicators of coastal conditions (Orth et al., 2006).
Estimating the abundance of seagrass is critical to judging seagrasses ecosystem functions (Hemminga and Duarte, 2000). Multiple methods are used to determine seagrass abundance, all evaluating one or more of the following categories: seagrass leaf cover, biomass, shoot density and canopy height (Short and Coles, 2001). Transects are specifically used to determine seagrass abundance over a range of water depths (Short and Coles, 2001).
Determining the maximum water depth at which seagrass grows is critical during EIA and environmental monitoring programs. This metric can be used as a proxy for estimating any anthroprogenic impact on the local marine ecosystem, of which impacts will be most significant to the maximum water depth at which seagrass grows.
There exist various species of seagrass, with Zostera marina or “eelgrass” being the specific seagrass species which is monitored as part of the Lynetteholm EIA, and therefore also the species of seagrass which models in this report have been trained on. However, in the context of using computer vision, seagrasses’ characteristic leaf morphology, density variations contributing to distinctive underwater meadow structures, and consistent color spectra suggest that methods applied to eelgrass can possibly work on other taxonomically similar species, and vice versa.
I-B Related Work
Multiple other studies have pursued similar efforts to automate eelgrass detection, classification and coverage estimation based on analysis of image data using various methods including DNNs. Prior work for seagrass coverage estimation includes segmenting images into superpixels and training a logistic regression classifier to distinguish image patches between seagrass and background (Reus et al., 2018). This method might not generalize well to images at different resolutions, scale of objects or water depths because of the fixed patch size constraint and the dependency on the quality of the superpixels. We propose a new approach based on estimating seagrass coverage efficiently from binary classifications in an image stream.
In SeaGrassDetect (Sengupta et al., 2020), the authors use unlabeled image data collected by a scuba diver towed by a boat to detect seagrass and estimate its coverage. They circumvent the necessity of labeled ground truth data by leveraging the structural characteristics of seagrass. They apply edge and line detection to extract the number of eelgrass stalks in an image and the total length of the lines which are combined with a weighted threshold strategy to infuse a Gaussian mixture model for binary classification. In our work we use an end-to-end seagrass detection approach where the entire image is fed into a DNN, removing the dependency on the sensitivity of feature engineering like line detection.
Previous studies incorporate the training of DNN models to detect a specific seagrass species called Halophila ovalis in images, captured in controlled lab environments with artificial lighting and real UW environments (Noman et al., 2023), (Moniruzzaman et al., 2019). Another line of research is to optimize DNNs for multi-species seagrass detection and classification (Raine et al., 2020) or for specific hardware constraints (Wang et al., 2020). We focus on the detection of a single seagrass species in the wild and investigate the performance of multiple DNN architectures on this task.
Other methods for seagrass monitoring include remote sensing (Veettil et al., 2020) collected from satellite images (Riegl and Purkis, 2005), aerial footage (Roelfsema et al., 2015), acoustic sensors (Gumusay et al., 2019), scuba divers (Sengupta et al., 2020) or remotely operated vehicles (Finkl and Makowski, 2016). However we focus on analyzing large amounts of unstructured video data gathered in the wild from EIA transect videos. This allows us to adhere to already established standard operating procedures for quality assurance in monitoring efforts as part of EIAs.
II Materials and Methods
II-A Data Acquisition
@@
Figure 2: (a) - (c) Sample images of areas with no, sparse and dense submerged aquatic vegetation (SAV). (d) A map of the Copenhagen harbor with transect lines (blue) and SAV data overlaid. All SAV maps are based on ©  Copernicus Sentinel-2 data from 2023-06-13. All SAV maps are based on analysis of Sentinel-2 imagery from 2018.
@@
We gather a novel annotated image data set of Zostera marina from six EIA transect surveys in and outside the Copenhagen harbour, covering an area of roughly $10\textrm{km}^{2}$. In total we consider 20 survey transects over two consecutive days in June 2023 during daytime. Each survey produces a transect video of the local benthic environment. During an eelgrass transect survey, an underwater camera is mounted on a sled and towed behind the survey vessel along a planned route of interest. The sled ensures the camera maintains a consistent angle and prevents it from swinging in the water current. A technician or biologist on the vessel monitors the depth of the sled and monitors its height above the seafloor, adjusting its height to avoid obstacles or to perform necessary maintenance such as cleaning the camera lens. After the field data collection, a marine biologist reviews the camera footage and adds annotations to a spreadsheet. These annotations include timestamps, the vessel’s location, and any notable changes in eelgrass coverage or other observations. All videos are filmed with an underwater camera with a 1/3” 2 Mega pixel Sony CMOS sensor.
@@
As part of the EIA of the Lynetteholm, project DHI performs 20 eelgrass video transects annually. Out of all 20 video transects of 2023 we choose a subset of six transect videos for their high degree in variability in visual content, i.e. changing seabed, as well as their roughly equal class distribution of eelgrass present vs. absent images. We estimate this based on earth observation data provided by DHI as part of the EIA, which showed SAV areas of interest (see Fig. 2(d)), and initial data inspection. Each selected transect video is between 5-60 minutes long and recorded at 30 FPS while the survey vessel sails at roughly 1m/s. We extract frames at 5 FPS from the survey videos in order to strike a balance between the volume of data and the informational content. This also ensures that consecutive frames are sufficiently distinct. In total we extracted 340,000 images from our selected transect videos. We crop all frames to 1280x500 pixels to remove any overlay information (see samples in Fig. 2). From each transect video we randomly sample images to create the dataset, resulting in total of 16,000 images.
Transect Name Length (km) Water Depth (m) Video Length (min)
LYT-5 2.00 $5.5\rightarrow 12.4$ 20
LYT-9 0.29 $2.7\rightarrow 7.4$ 5
LYT-10 0.45 $3.0\rightarrow 6.0$ 10
LYT-12 0.50 $3.4\rightarrow 12.8$ 10
LYT-14 3.00 $13.8\rightarrow 14.1$ 60
LYT-20 1.20 $9.5\rightarrow 9.6$ 45
@@
TABLE I: Transect Details. All transects were filmed in June 2023 over two consecutive days during daytime.
II-B Data Annotation and Platform

### v2

@@ Abstract
Seagrass meadows play a crucial role in marine ecosystems, providing benefits such as carbon sequestration, water quality improvement, and habitat provision. Monitoring the distribution and abundance of seagrass is essential for environmental impact assessments and conservation efforts. However, the current manual methods of analyzing underwater video data to assess seagrass coverage are time-consuming and subjective. This work explores the use of deep learning models to automate the process of seagrass detection and coverage estimation from underwater video data. We create a new dataset of over 8,300 annotated underwater images, and subsequently evaluate several deep learning architectures, including ResNet, InceptionNetV3, DenseNet, and Vision Transformer for the task of binary classification on the presence and absence of seagrass by transfer learning. The results demonstrate that deep learning models, particularly Vision Transformers, can achieve high performance in predicting eelgrass presence, with AUROC scores exceeding 0.95 on the final test dataset. The application of underwater image enhancement further improved the models’ prediction capabilities. Furthermore, we introduce a novel approach for estimating seagrass coverage from video data, showing promising preliminary results that align with expert manual labels, and indicating potential for consistent and scalable monitoring. The proposed methodology allows for the efficient processing of large volumes of video data, enabling the acquisition of much more detailed information on seagrass distributions in comparison to current manual methods. This information is crucial for environmental impact assessments and monitoring programs, as seagrasses are important indicators of coastal ecosystem health. This project demonstrates the value that deep learning can bring to the field of marine ecology and environmental monitoring.
keywords:
[…]
{graphicalabstract}[Uncaptioned image]
{highlights}
Deep neural networks can effectively automate seagrass presence detection in underwater videos from environmental impact assessments, reducing reliance on manual annotations.
Transfer learning combined with underwater image enhancement techniques improves deep learning capabilities for seagrass detection tasks.
Our custom data annotation platforms support labeling of large amounts of data collected from environmental impact assessments.
Our methodology serves as a supportive tool for marine biologists, enabling more consistent, scalable, and objective environmental monitoring.
1 Introduction
Seagrasses are a type of submerged aquatic vegetation (SAV) found in shallow waters alongside every continent in the world except Antarctica (Novak and Short, 2020). As a marine flowering plant, they are of fundamental importance to the health and integrity of coastal ocean habitats, biodiversity, global climate, and human economic stability as well as global food security (Unsworth et al., 2019a; Cullen-Unsworth et al., 2014). Seagrass meadows are monitored worldwide as part of environmental impact assessments (EIAs) and environmental monitoring programs because they provide important ecosystem services, such as carbon sequestration, water quality improvement, providing food and habitat, and acting as a biological indicator of local ecosystems (Novak and Short, 2020). Seagrass meadows serve as vital hubs of biodiversity, offering nourishment and shelter to a wide range of organisms, including endangered species like dugongs and commercially valuable ones such as fish and shrimp, as well as microbes and invertebrates (Fry and Parker, 1979). They are estimated to support up to 20% of global fishery productivity (Unsworth et al., 2019b), making them key contributors to marine food webs and fisheries management. Although seagrasses cover only 2% of the ocean’s seabed, they account for over 15% of the carbon annually sequestered in marine environments (Fourqurean et al., 2012). Despite ongoing conservation efforts, seagrass meadows are declining globally at an alarming rate due to human activities and coastal development, which is further exacerbated by the effects of climate change (Waycott et al., 2009). These factors collectively contribute to a stark decline in water quality, and threaten the long-term survival of seagrass habitats. Due to seagrasses immediate vulnerability to environmental factors and their key ecosystem functions, they are important biological indicators and ecological status of coastal conditions (Orth et al., 2006; Balsby et al., 2013).
Estimating the abundance of seagrass meadows is critical to judging seagrasses’ ecosystem functions (Hemminga and Duarte, 2000). The fundamental challenge here is to acquire suitable data in a structured, automated and efficient manner (Unsworth et al., 2019a). Methods for seagrass monitoring include analysis of images from remote sensing (Veettil et al., 2020), for example collected from satellite (Riegl and Purkis, 2005), and from aerial footage (Roelfsema et al., 2015). Yet these methods fail to distinguish between different types of seagrass and lack fine-grained resolution. Additionally, aerial surveys are heavily limited by water turbidity and favorable weather conditions. Traditionally, scuba divers have been deployed to conduct seagrass monitoring (Sengupta et al., 2020), however this is extremely restricted due to the required amount of human labor. There is a trend towards monitoring with autonomous underwater vehicles which adds new challenges and possibilities to automated seabed mapping (Moniruzzaman et al., 2019a).
Environmental Impact Assessments (EIAs) conducted offshore are intended to monitor and manage the environmental impacts of coastal developments such as offshore construction sites in order to ensure environmental compliance with different stakeholders across the marine infrastructure sector. During EIAs, monitoring is often conducted in the form of video transects, among other methods, to determine the seagrass abundance over a range of water depths (Short and Coles, 2001). To achieve this, a camera is mounted onto a sled and towed behind a vessel to capture underwater video footage. The recorded videos are then annotated on the presence or absence of seagrass by marine biologists. This method provides several advantages; a large spatial coverage, more fine-grained image data of the underwater environment, and the possibility of repeated monitoring in a structured and efficient manner. Similarly to other image-based monitoring methods, the subsequent data labeling task is extremely tedious as it is done manually and thus requires a vast amount of working hours.
@@ 1 Introduction
Figure 1: The SeagrassFinder Project Pipeline: (1) a vessel performs transects surveys by towing a sled with a camera along the seabed and (2) records videos of the underwater environment. The extracted images (3) are then labeled by human annotators (4) and used to train deep neural networks (DNNs) on the task of detecting eelgrass (5). A trained DNN can now be used to replace the time-consuming manual annotation process.
With the advent of deep learning, our capacity to analyze and process large amounts of data has been transformed significantly. Specifically deep neural networks (DNNs) have shown remarkable success in a wide range of computer vision tasks such as image classification, semantic segmentation and object detection. DNNs achieve this by learning information directly from raw pixel input data without the need for manual feature engineering. Especially convolutional neural networks (Krizhevsky et al., 2012) have massively improved our ability to automatically extract relevant features and have become the new standard for many computer vision tasks. DNN training and prediction for the task of seagrass monitoring is a promising direction, because DNNs can conduct image classifications in an automated, repeatable and efficient manner and offer generalizability across a variety of image qualities suffered from underwater images, such as blur, color-distortion or low contrast and low light. Meanwhile, seagrass stalks can have complex structures and shapes, thus, even single-species seagrass detection is a hard challenge. Moniruzzaman et al. (Moniruzzaman et al., 2019b) detect single seagrass stalks of Halophila ovalis with a Faster R-CNN model on their own compiled data set ECUHO, containing images in a laboratory setting and images taken in the wild. Noman et al. (Noman et al., 2023) achieve a higher mAP performance of 0.484 on the ECUHO data set, with two popular object detection models, YOLOv5 and EfficientNetD7. Sengupta et al. (Sengupta et al., 2020) also investigate single-species detection. They use traditional feature extraction methods in combination with a Gaussian mixture model for binary classification on absence or presence of seagrass.
Multi-species seagrass classification is often evaluated on the patch-based DeepSeagrass data set (Raine et al., 2020), on which the authors achieve an accuracy of 92.4% via semi-automatically patch labeling and a convolutional DNN. In the work by Noman et al. (Noman et al., 2021), the authors achieve overall higher accuracy with a semi-supervised learning approach. Latest advances by unsupervised curriculum learning by Abid et al. (Abid et al., 2024) perform best at 90.12% precision.
Prior work on image segmentation includes feature extraction in combination with logistic model trees (Bonin-Font et al., 2016), superpixel-based solutions (Reus et al., 2018), later revised by Weidmann et al. (Weidmann et al., 2019), and patch-based approaches (Burguera, 2020), reaching 95.5% of accuracy. Further work on a semantic segmentation on a pixel basis was conducted first with a VGG16 encoder and FCN8 decoder (Martin-Abadal et al., 2018), later with an attention-based encoder-decoder modules (Wang et al., 2020b), as well as with an encoder similar to MobileNet (Wang et al., 2020a). Most recently, Raine et al. (Raine et al., 2024) have combined large language models as a supervisory signal with unsupervised contrastive pre-training for coarse seagrass segmentation.
@@
We address the challenge of estimating seagrass presence and coverage in the wild from vast amounts of raw, unstructured and unfiltered video data taken in realistic environmental conditions. We achieve this by deploying a custom-made annotation platform for efficient manual labeling, as well as training state-of-the-art deep learning models for classifying whether eelgrass is present or absent in an image. To the best of our knowledge we are the first to deploy Vision Transformers (Dosovitskiy, 2020) to detect the presence of eelgrass in underwater images. In this work, we emphasize the distinction between spatial seagrass coverage - referring to the actual vegetation density - and pixel-based coverage, which determines the proportion of image pixels classified as seagrass. This is often overlooked in current contributions, yet we argue spatial seagrass coverage holds greater ecological relevance, as it reflects actual vegetation density rather than just pixel-level detection. Specifically we contribute the following:
@@
    a strategy for effectively streamlining human labeling efforts of human experts and non-experts on underwater seagrass images and compiling the annotated data set into an efficient deep learning data pipeline (see Fig. 1),
@@
    an analysis of state-of-the-art DNN models and their ability to classify a range of challenging underwater images according to the presence of eelgrass using a transfer learning mechanism, as well as an investigation of the effect of UW image enhancement on DNN performance,
@@
    a new experimental approach for further utilizing the above-mentioned DNN models for estimating visual eelgrass coverage for an image stream from binary predictions as a post-processing step,
@@
    a pipeline which dramatically reduces the time and financial resources required for human annotation, and paves the way for automated environmental monitoring for example by means of autonomous underwater vehicles.
The remainder of this paper is structured as follows: Sect. 2 describes the data strategy, our deep learning methodology for eelgrass detection, as well as our novel eelgrass coverage estimation. We present our experimental design in Sect. 3 and our results in section Sect. 4. Finally, we present our discussion and conclusion in Sect. 5 and Sect. 6.
2 Materials and Methods
In this section we present our data strategy, followed by our deep learning methodology on the task of eelgrass detection and finally our novel eelgrass coverage estimation as a postprocessing step.
2.1 Data Acquisition
@@
Figure 3: (a) - (c) Sample images of areas with no, sparse and dense submerged aquatic vegetation (SAV). (d) A map of the Copenhagen harbor with transect survey lines (blue) and SAV data overlaid. All SAV maps are based on ©  Copernicus Sentinel-2 data from 2023-06-13. All SAV maps are based on analysis of Sentinel-2 imagery from 2018 (map created with QGIS v.3.30.2).
@@
The dataset used in this project originates from DHI’s Environmental Impact Assessment (EIA) and baseline reporting of the Lynetteholm project, a new artificial peninsula being created to extend the coastline of northern Copenhagen in Denmark. We focus on one species of seagrass, namely Zostera marina, commonly known as “eelgrass” which is the target specific seagrass species monitored in the Lynetteholm EIA (see samples in figures 3(a), 3(b) and 3(c). We compile a novel annotated image dataset of Zostera marina from six transect surveys in- and outside the Copenhagen harbor, covering an area of roughly $10\textrm{km}^{2}$. In total we consider 20 transect surveys (marked blue in Fig. 3(d)) over two consecutive days in June 2023 during daytime. Each transect survey produces a transect video of the local benthic environment, the lowest level in a body of water, close to the seabed. During a transect survey, an underwater camera is mounted on a sled, a mechanical device, designed to hover at a roughly consistent height above the seabed while towed behind a survey vessel along a planned route of interest. The sled ensures the camera maintains a consistent angle and prevents it from swinging in the water current. A technician or biologist on the vessel monitors the depth of the sled and monitors its height above the seafloor, adjusting its height to avoid obstacles or to perform necessary maintenance such as cleaning the camera lens. After the field data collection, a marine biologist reviews the camera footage and adds annotations on seagrass abundance to a spreadsheet. These annotations include timestamps, the vessel’s location, and any notable changes in eelgrass coverage or other observations. All videos are filmed with an underwater camera with a 1/3” 2 Mega pixel Sony CMOS sensor.
@@
As part of the EIA of the Lynetteholm project, DHI performs 20 transect surveys annually, each producing one transect video. Out of all 20 transect videos of 2023, we choose a subset of six transect videos, based on two factors. We are primarily interested in transect videos where the visual content shows a high degree in variability. We speculate this enhances a DNN’s capability for generalization and classification accuracy, for example when confronted with changes in seabed textures. We additionally choose transect videos for their roughly equal class distribution of eelgrass present vs. absent images, in order to avoid model bias in the subsequent DNN training. A balanced data set ensures that a DNN learns both classes well, without favoring a more frequently represented class. We estimate this based on earth observation data provided by DHI as part of the EIA, which showed Submerged Aquatic Vegetation (SAV) areas of interest (see Fig. 3(d)), and initial data inspection. Each selected transect video is between 5-60 minutes long and recorded at 30 FPS while the survey vessel sails at roughly 1m/s. We extract frames at 5 FPS from the survey videos in order to strike a balance between the volume of data and the informational content. This ensures that consecutive frames are sufficiently distinct. In total we extracted 340,000 images from our selected transect videos. We crop all frames to 1280x500 pixels to remove any overlay information (see examples in Fig. 3). From each transect video we randomly sample images to create the dataset, resulting in a total of 16,000 images.
@@
Table 1: Transect Details. All transects were filmed in June 2023 over two consecutive days during daytime.
Transect Name Length (km) Min. Water Depth (m) Max. Water Depth (m) Video (min)
LYT-5 2.00 5.5 12.4 20
LYT-9 0.29 2.7 7.4 5
LYT-10 0.45 3.1 6.0 10
LYT-12 0.50 3.4 12.8 10
LYT-14 3.00 13.8 14.1 60
LYT-20 1.20 9.5 9.6 45
2.2 Data Annotation and Platform